OpenStudy (anonymous):

write a recursive formula for 8,6,4,2,0 then find a14

OpenStudy (polaris_s0i):

you're recursion is subtraction by 2. \[ y_{n+1} = y_n - 2 \] this then gives you: \[ y_1 = 8 - 2 = 6 \] \[ y_2 = 6 - 2 = 4 \] however you can "solve" this: \[y_1 = (y_0 - 2)\] \[y_2 = ((y_0 -2) -2) = y_0 - 4 = y_0 - 2*2\] \[y_3 = (((y_0 - 2) - 2) - 2) = y_0 -6 = y_0 - 2*3\] so in general you get: \[ y(n) = y_0 - 2n\] so then at 14 you would get: \[y(14) = 8 - 2*14 = -20\]